The charge (C) through a resistor is given by the equation: q=4t^3-9t^2. What is the instantaneous current through the resistor at t=5.0s. (i=dq/dt)
Answer by Alan3354(69443)

The charge (C) through a resistor is given by the equation: q=4t^3-9t^2. What is the instantaneous current through the resistor at t=5.0s.
----------
q' = dq/dt = 12t^2 - 18t
q'(5) = 12*25 - 18*5 = 210 --- units not given.